# Height filtrations and base loci on flag bundles over a curve

### v1

@@ 2.4. Height function induced by $(F/P,\mathcal{L}_{\lambda}$)
2.5. A height lower bound in Schubert cells
For $w\in W_{Q}\backslash W/W_{P}$, write $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P$, $\mathcal{X}_{w}=F_{Q}\times_{Q}\overline{QwP}/P$, $C_{w}=\mathcal{C}_{w,K}$ and $X_{w}=\mathcal{X}_{w,K}$. We shall show that
Proposition 2.4.
For any $x\in C_{w}(\overline{K})$, $h_{\mathcal{L}_{\lambda}}(x)\geq\langle{\mathrm{deg}}F_{Q},w\lambda\rangle$.
@@ 2.5. A height lower bound in Schubert cells
$F_{P}\hookrightarrow F_{P}\times_{P}G\cong F,\quad F_{Q}\hookrightarrow F_{Q}\times_{Q}G\cong F.$
@@
Now we can prove Proposition 2.4.
@@
Let $L/K$ be any finite field extension. Assume that $x\in C_{w}(L)$. Consider the following commutative diagram
@@
2.6. Height filtration and successive minima of $h_{\mathcal{L}_{\lambda}}$
In this subsection, we employ Ballaÿ’s theorem [2, Theorem 1.2] to compute essential minimum. We start by finding the maximal slope of $\pi_{*}\mathcal{L}_{\lambda}|_{\mathcal{X}_{w}}$.
@@ 2.6. Height filtration and successive minima of $h_{\mathcal{L}_{\lambda}}$
Schieder computed the Harder-Narasimhan filtration of $F_{Q}\times_{Q}H^{0}(\overline{QwP}/P,M_{\lambda})$ in [23, §5.1]. Let $V/k$ be a finite dimensional $Q$-representation with weight decomposition $V=\bigoplus_{\nu\in X(T)}V[\nu]$. The $Q$-filtration $V_{\bullet,{\mathrm{deg}}(F_{Q})}$ of $V$ is defined by
$V_{t,{\mathrm{deg}}(F_{Q})}=\bigoplus_{\langle{\mathrm{deg}}(F_{Q}),\nu\rangle\geq t}V[\nu].$
@@
When $Q=B$, $H^{0}(\overline{BwP}/P,M_{\lambda})$ has a single highest weight $w\lambda$ by the theory of Demazure modules [17, Chapter 14]. For general $Q$, let $\pi:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ be the canoniacl projection. Note that
@@
Corollary 2.10.
@@
The essential minimum $\zeta_{1}(h_{\mathcal{L}_{\lambda}},X_{w})$ of $h_{\mathcal{L}_{\lambda}}$ on $X_{w}$ is $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$.
@@
By Ballaÿ’s theorem [2, Theorem 1.2],
$\quad\quad\quad\zeta_{1}(h_{\mathcal{L}_{\lambda}},X_{w})$ $=\displaystyle\lim_{n\rightarrow\infty}\frac{\mu_{\max}(\pi_{*}\mathcal{L}_{n\lambda}|_{\mathcal{X}_{w}})}{n}$
Theorem 2.11.
The height filtration of $h_{\mathcal{L}_{\lambda}}$ on $X$ is given by successively deleting Schubert cells $C_{w}$ for $w\in W_{Q}\backslash W/W_{P}$, i.e.
In particular, successive minima are $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ and Zhang’s successive minima are $e_{i}=\min\big{\{}\zeta_{w}:\ell(w)=d-i+1\big{\}}$.
Proof.
On $X_{w}$, $\zeta_{1}(h_{\mathcal{L}_{\lambda}},X_{w})=\zeta_{w}$ and for any $x\in C_{w}$, $h(x)\geq\zeta_{w}$. Thus $Z_{\zeta_{w}}(X_{w})\subseteq X_{w}\backslash C_{w}=\bigcup_{w^{\prime}<w}X_{w^{\prime}}$. For a closed subvariety $Y$ of $X$, let $Z_{t}(Y)$ be the Zariski closure in $Y$ of $\big{\{}y\in Y:h(y)<t\big{\}}$. We claim that $Z_{\zeta_{w}}(X_{w})=X_{w}\backslash C_{w}=\bigcup_{w^{\prime}<w}X_{w^{\prime}}$.
In fact, suppose conversely that $Z_{\zeta_{w}}(X_{w})\subsetneq X_{w}\backslash C_{w}$. Then there exists $X_{w^{\prime}}\subsetneq X_{w}$ such that $Z_{\zeta_{w}}(X_{w})\cap X_{w^{\prime}}\subsetneq X_{w^{\prime}}$. This forces $Z_{\zeta_{w}}(X_{w^{\prime}})\subseteq Z_{\zeta_{w}}(X_{w})\cap X_{w^{\prime}}\subsetneq X_{w^{\prime}}$. But on $X_{w^{\prime}}$, the essential minimum is $\zeta_{w^{\prime}}<\zeta_{w}$, so $Z_{\zeta_{w}}(X_{w^{\prime}})=X_{w^{\prime}}$. We get contradiction and we have thus proved that on $X_{w}$, the set $Z_{t}(X_{w})=X_{w}$ when $t>\zeta_{w}$ and $Z_{t}(X_{w})=X_{w}\backslash C_{w}$ when $t\leq\zeta_{w}$.
3. Reinforcement of Zhang’s inequality into an equality
In this section, we shall establish the following Zhang-type equality on flag varieties over function fields: the height of the variety in fact equals to a weighted average of the successive minima.
Theorem 3.1.
Let $q:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ be the canoniacl projection. For $w\in W_{Q}\backslash W/W_{P}$, define $m_{w}=\#q^{-1}(w)$. Then
$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\#W/W_{P}}\sum_{w\in W_{Q}\backslash W/W_{P}}m_{w}\zeta_{w}.$
We shall establish this theorem by convex geometry. Let $E$ be a vector bundle on $C$ with HN filtration $0=E_{0}\subseteq E_{1}\subseteq\cdots\subseteq E_{r}=E$ with successive slopes $\mu_{i}=\mu(E_{i}/E_{i-1})$. Chen defines a measure $\nu_{E}$ on $\mathbb{R}$ by
$\nu_{E}=\frac{1}{\operatorname{rank}(E)}\sum_{i=1}^{r}\operatorname{rank}(E_{i}/E_{i-1})\delta_{\mu_{i}},$
where $\delta_{a}$ is the Dirac measure on $\mathbb{R}$ supported at $a\in\mathbb{R}$. For any $\varepsilon>0$ and any measure $\mu$, let $T_{\varepsilon}\mu$ be the unique measure such that $\int_{\mathbb{R}}f(x)\mathrm{d}T_{\varepsilon}\mu(x)=\int_{\mathbb{R}}f(\varepsilon x)\mathrm{d}\mu(x)$ for any compactly-supported continuous function $f:\mathbb{R}\longrightarrow\mathbb{R}$.
Theorem 3.2 (Chen; [11], Theorem 1.1).
Let $\pi:\mathcal{X}\longrightarrow C$ be a projective flat morphism and $\mathcal{L}$ be a line bundle on $\mathcal{X}$. Assume $\mathcal{L}$ is generically big. Then
  1. (1)
    $\big{\{}T_{\frac{1}{n}}\nu_{\pi_{*}\mathcal{L}^{n}}\big{\}}_{n}$ converges weakly. We denote the limit measure by $\nu^{\pi}_{\mathcal{L}}$.
  2. (2)
    $\operatorname{Vol}(\mathcal{L})=\dim\mathcal{X}\cdot\operatorname{Vol}(\mathcal{L}_{K})\cdot\int_{\mathbb{R}}x_{+}\;\mathrm{d}\nu^{\pi}_{\mathcal{L}}$, where $x_{+}=\max\{0,x\}$.
Let $w_{0}\in W$ be the longest element and fix a reduced decomposition $w_{0}=s_{1}\cdot s_{2}\cdots s_{{\ell(w_{0})}}$. Recall that $\lambda:P\longrightarrow\mathbb{G}_{m}$ is a strictly anti-dominant character and $N=\dim G/P$. Let $\Delta_{\lambda}=\Delta_{\underline{w}_{0},\lambda}\subseteq\mathbb{R}^{N}$ be the string polytope associated to $\lambda$ with respect to $\underline{w}_{0}=(s_{1},\cdots,s_{\ell(w_{0})})$. It is equipped with a weight map $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$ where $\mathcal{P}_{\lambda}\subseteq X(T)_{\mathbb{R}}$ is the weight polytope associated to $\lambda$ (see [1, §1] for details). The following properties will be useful:
  1. (1)
    The volume $\operatorname{Vol}(\Delta_{\lambda})$ of the convex polytope equals to the volume $\operatorname{Vol}(M_{\lambda})$ of the line bundle.
  2. (2)
    Integral points $\Delta_{\lambda}\cap\mathbb{Z}^{N}$ parametrize a basis of $H^{0}(G/P,M_{\lambda})$. In particular, #$\#\Delta_{\lambda}\cap\mathbb{Z}^{N}=\dim H^{0}(G/P,M_{\lambda})$.
  3. (3)
    Lattice points $\mathcal{P}_{\lambda}\cap X(T)$ parameterize weights occuring in $H^{0}(G/P,M_{\lambda})$ and $p$ sends an integral point to the underlying weight of the corresponding basis vector.
  4. (4)
    $\Delta_{\lambda_{1}}+\Delta_{\lambda_{2}}=\Delta_{\lambda_{1}+\lambda_{2}}$ and $p_{\lambda_{1}}+p_{\lambda_{2}}=p_{\lambda_{1}+\lambda_{2}}$. Here the sum is the Minkowski sum. In particular, the diagram
    ${\Delta_{\lambda}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p_{\lambda}}$$\scriptstyle{\cdot n}$${\Delta_{n\lambda}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p_{n\lambda}}$${\mathcal{P}_{\lambda}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\cdot n}$${\mathcal{P}_{n\lambda}}$
    is commutative for any $n\in\mathbb{N}$, where $\cdot n$ is the scalar multiplication by $n$.
@@ 3. Reinforcement of Zhang’s inequality into an equality
Let $\mu$ be the Lebesgue measure on $\mathbb{R}^{N}$ normalized with respect to $\mathbb{Z}^{N}$.
Lemma 3.3.
The following equalities hold:
$\nu_{\pi_{*}\mathcal{L}^{n}_{\lambda}}=\frac{1}{\#(\Delta_{n\lambda}\cap\mathbb{Z}^{N})}\sum_{x\in\Delta_{n\lambda}\cap\mathbb{Z}^{N}}\delta_{{\mathrm{deg}}F_{Q}\circ p(x)},$
$T_{\frac{1}{n}}\nu_{\pi_{*}\mathcal{L}_{\lambda}^{n}}=\frac{1}{\#(\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N})}\sum_{x\in\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}\delta_{{\mathrm{deg}}F_{Q}\circ p(x)},$
@@
$\nu^{\pi}_{\mathcal{L}_{\lambda}}=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\cdot({\mathrm{deg}}F_{Q}\circ p)_{*}\mu.$
Proof.
Bearing Lemma 2.8, Proposition 2.9 and the aforementioned properties of string polytopes at hand, we can carry over the argument in [11, Proposition 3.5] to show the first equality. Appling $T_{1/n}$ and taking limits, we obtain the second and third equalities. ∎
Theorem 3.4.
The variety height $h_{\mathcal{L}_{\lambda}}(X)$ of $X$ is $\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}{\mathrm{deg}}(F_{Q})\circ p\;\mathrm{d}\mu$.
Proof.
We may consider $\mathcal{L}_{\lambda}+tf$ for large $t$ so that $\mathcal{L}_{\lambda}+tf$ is ample, where $f$ is the line bundle given by a vertical fiber. The measure $\mu^{\pi}_{\mathcal{L}_{\lambda}+tf}$ is supported on $\mathbb{R}_{+}$ and is a right shift of $\mu^{\pi}_{\mathcal{L}_{\lambda}}$ by $t$, so Theorem 3.2 says
$\quad\quad\quad(\mathcal{L}_{\lambda}+tf)^{N+1}$ $=(N+1)\operatorname{Vol}(\mathcal{L}_{\lambda,K})\int_{\mathbb{R}}x\;\mathrm{d}\nu^{\pi}_{\mathcal{L}_{\lambda}+tf}$
$=(N+1)\int_{\Delta_{\lambda}}{\mathrm{deg}}(F_{Q})\circ p+t\;\mathrm{d}\mu$
where the second equality is due to Lemma 3.3 and $\operatorname{Vol}(\mathcal{L}_{\lambda,K})=\operatorname{Vol}(\Delta_{\lambda})$. Thus
$h_{\mathcal{L}_{\lambda}+tf}(X)=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}{\mathrm{deg}}(F_{Q})\circ p\;\mathrm{d}\mu+t.$
Note that $h_{\mathcal{L}_{\lambda}+tf}(X)=h_{\mathcal{L}_{\lambda}}(X)+t$. So substracting $t$, we get
$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}{\mathrm{deg}}(F_{Q})\circ p\;\mathrm{d}\mu.$
To prove Theorem 3.1, we need the following general result.
@@
For any linear function $f:X(T)_{\mathbb{R}}\longrightarrow\mathbb{R}$,
@@
We start with two lemmata.
@@
Note that the Lebesgue measure $\mu$ on $\Delta_{\lambda}$ is the limit of the Dirac measure $\delta_{\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}$, i.e. $\mu=\lim_{n}\delta_{\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}$. Consider the map $\Delta_{n\lambda}\longrightarrow\mathcal{P}_{n\lambda}$ and note that it sends integral points to its corresponding weights. Dividing by n, we see that
@@
$\operatorname{mult}_{n}:\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)\longrightarrow\mathbb{N},\quad\mu\longmapsto\dim H^{0}(G/P,L_{n\lambda})[n\mu].$
@@
$H^{0}(G/P,L_{\lambda})$ is a simple $G$-module so all weights $\mu$ occuring in $H^{0}(G/P,L_{\lambda})$ sharing the same restriction to $Z(G)$. So by the above claim, the lemma is true for lattice points $\mathcal{P}_{\lambda}\cap X(T)$. Note that the map $\mu\longmapsto\sum_{w\in W}w\mu$ is linear, we are done. ∎
@@
Proof of Theorem 3.1.
Take $f={\mathrm{deg}}(F_{Q})$ in Proposition 3.5 and apply Theorem 3.4. ∎
Remark 3.8.
The fact that the variety height and successive minima are given by
$\displaystyle h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\operatorname{Vol(\Delta_{\lambda})}}\int_{\mathcal{P}_{\lambda}}{\mathrm{deg}}(F_{Q})\;\mathrm{d}p_{*}\mu,\quad\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$
shows that the Arakelov geometry of $(F/P,\mathcal{L}_{\lambda})$ depends only on the function ${\mathrm{deg}}(F_{Q})$ on $X(T)_{\mathbb{Q}}$, which might be surprising at the first glance.
This is in fact because the Arakelov geometry of $(F/P,\mathcal{L}_{\lambda})$ is determined by the Harder-Narasimhan filtration of $\pi_{*}\mathcal{L}_{n\lambda}=F_{Q}\times_{Q}H^{0}(G/P,M_{n\lambda})$ (Lemma 2.8) for all $n\geq 1$, and the unipotent part of $Q$ acts trivially [23, Lemma 5.1].

### v2

@@ 2.5. Arakelov geometry on flag varieties over function fields
$F_{P}\hookrightarrow F_{P}\times_{P}G\simeq F,\quad F_{Q}\hookrightarrow F_{Q}\times_{Q}G\simeq F.$
@@
Proof.
This is written in the proof of [23, Theorem 4.1]. See [23, §4.5.3]. ∎
@@
Let $L$ be any finite field extension of $K$. Assume that $x\in C_{w}(L)$. Consider the following commutative diagram
@@
2.7. Proof of Proposition 2.5
In this subsection, we employ Ballaÿ’s theorem [1, Theorem 1.2] to compute the essential minimum. Let $\pi:F/P\longrightarrow C$ be the structure morphism. We start by finding the maximal slope of $\pi_{*}\mathcal{L}_{\lambda}|_{\mathcal{X}_{w}}$.
@@
Schieder computed the Harder-Narasimhan filtration of $F_{Q}\times_{Q}H^{0}(\overline{QwP}/P,M_{\lambda})$ in [23, §5.1]. To state his result, we fix the following notations which will be used frequently.
Definition 2.11.
Let $V$ be a finite dimensional $Q$-representation over $k$ with weight decomposition $V=\bigoplus_{\mu\in X(T)}V[\mu]$. The $Q$-filtration $V_{\bullet,{\mathrm{deg}}(F_{Q})}$ of $V$ is defined by
$V_{t,{\mathrm{deg}}(F_{Q})}=\bigoplus_{\langle{\mathrm{deg}}(F_{Q}),\mu\rangle\geq t}V[\mu].$
@@
When $Q=B$, $H^{0}(\overline{BwP}/P,M_{\lambda})$ has a single highest weight $w\lambda$ by the theory of Demazure modules [14, Chapter 14]. For general $Q$, let $\pi:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ be the canoniacl projection. Note that
@@
Proof of Proposition 2.5.
By Ballaÿ’s theorem [1, Theorem 1.2],
$\quad\quad\quad\zeta_{\operatorname{ess}}(X_{w})$ $=\displaystyle\lim_{n\rightarrow\infty}\frac{\mu_{\max}(\pi_{*}\mathcal{L}_{n\lambda}|_{\mathcal{X}_{w}})}{n}$
3. String polytopes and the Boucksom-Chen concave transform
Keep the notations in §2. Let $w_{0}\in W$ be the longest Weyl element and fix a reduced decomposition $\underline{w_{0}}=s_{i_{1}}s_{i_{2}}\cdots s_{i_{N}}$, $N=\dim G/B$. Then for the antidominant weight $\lambda:P\longrightarrow\mathbb{G}_{m}$, we have the string valuation $\nu$ and the string polytope $\Delta_{\lambda}\subseteq\mathbb{R}^{N}$, which is a polytope of dimension $d=\dim G/P$. Let $\mathcal{P}_{\lambda}$ be the weight polytope and $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$ be the weight map.
The main theorem of this section is the following:
Theorem 3.1.
The Boucksom-Chen concave transform of ${\mathcal{L}_{\lambda}}$ (with respect to a specific valuation) is $\Delta_{\lambda}\longrightarrow\mathbb{R},x\longmapsto\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle$.
Let $\mu$ be the Lebesgue measure on the real span of $\Delta_{\lambda}$, normalized with respect to $\mathbb{Z}^{N}$. As a direct corollary, we have by [6] that
@@ 3. String polytopes and the Boucksom-Chen concave transform
$h_{\mathcal{L}_{\lambda}}(X)=\frac{\mathcal{L}_{\lambda}^{\dim\mathcal{X}}}{\dim\mathcal{X}\cdot L^{\dim X}}=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle\,\mathrm{d}\mu(x).$
The integration can be computed via Weyl symmetry:
Theorem 3.2.
Let $q:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ be the canoniacl projection. For $w\in W_{Q}\backslash W/W_{P}$, let $m_{w}=\#q^{-1}(w)$. Then
$h_{\mathcal{L}_{\lambda}}(X)=\frac{1}{\#W/W_{P}}\sum_{w\in W_{Q}\backslash W/W_{P}}m_{w}\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle.$
Note that $\zeta_{w}=\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ are successive minima of the height function $h_{\mathcal{L}_{\lambda}}$ on $X$ by Theorem 2.1. Thus this theorem builds an equality between the variety height and a weighted average of the successve mimina. One may recognize this as a refinement of Zhang’s inequality of successive minima.
3.1. Okounkov bodies and the Boucksom-Chen concave transform
Let $\mathbf{k}$ be a field and $Y$ be a normal projective variety of dimension $d$ over $\mathbf{k}$. Let $L$ be a big line bundle on $Y$, $V_{n}=H^{0}(Y,nL)$ and $V_{\bullet}=\bigoplus_{n\geq 0}V_{n}$. Let $\nu:V_{\bullet}\backslash\{0\}\longrightarrow\mathbb{Z}^{r}$ be a valuation with one dimensional leaves, i.e.
  • for $a\in\mathbf{k}^{\times}$ and $s\in V_{\bullet}\backslash\{0\}$, we have $\nu(as)=\nu(s)$.
  • for $s_{1},s_{2}\in V_{\bullet}\backslash\{0\}$, we have $\nu(s_{1}s_{2})=\nu(s_{1})+\nu(s_{2})$ and $\nu(s_{1}+s_{2})\geq\min\big{\{}\nu(s_{1}),\nu(s_{2})\big{\}}$.
  • the quotient space $\big{\{}s\in H^{0}(Y,nL):\nu(s)\geq a\big{\}}\Big{/}\big{\{}s\in H^{0}(Y,nL):\nu(s)>a\big{\}}$ is at most one dimensional.
Let $W_{\bullet}\subseteq V_{\bullet}$ be a graded subalgebra. We define $\Delta_{\nu}(W_{\bullet})\subseteq\mathbb{R}^{r}$ to be the closure of the set $\bigcup_{n\geq 1}\tfrac{1}{n}\nu(W_{n}\backslash\{0\})$. In particular, we define the Okounkov body $\Delta_{\nu}(Y,L):=\Delta_{\nu}(V_{\bullet})$.
A fundamental theorem is $\operatorname{Vol}(L)=d!\operatorname{Vol}(\Delta_{\nu}(Y,L))$ [16, Theorem 1.15], where the left hand side is the volume of the line bundle $L$ and the right hand side is the volume of the polytope $\Delta_{\nu}(Y,L)$.
Let $\big{\{}V_{\bullet}^{t}:t\in\mathbb{R}\big{\}}$ be a decreasing $\mathbb{R}$-filtration of $V_{\bullet}$ that is multiplicative, i.e. $V_{n_{1}}^{t_{1}}\cdot V_{n_{2}}^{t_{2}}\subseteq V_{n_{1}+n_{2}}^{t_{1}+t_{2}}$. The Boucksom-Chen concave transform associated to $\left(V_{\bullet},\nu,\big{\{}V_{\bullet}^{t}\big{\}}\right)$ is the function $G:\Delta_{\nu}(V_{\bullet})\longrightarrow\mathbb{R}\cup\{-\infty\}$, $x\longmapsto\sup\big{\{}t\in\mathbb{R}:x\in\Delta_{\nu}(V_{\bullet}^{t})\big{\}}$ [6]. $G$ is concave, as indicated in the name. In particular, $G$ is continuous in the interior.
3.2. String polytopes and weight polytopes
Let $w_{0}\in W$ be the longest Weyl element and fix a reduced decomposition $\underline{w_{0}}=s_{i_{1}}s_{i_{2}}\cdots s_{i_{N}}$, $N=\dim G/B$. In [19, 3], the authors constructed the string valuation $\nu=\nu_{\underline{w_{0}}}:H^{0}(G/P,nM_{\lambda})\backslash\{0\}\longrightarrow\mathbb{Z}^{N}$. The image $\nu(H^{0}(G/P,nM_{\lambda})\backslash\{0\})=\big{\{}v_{1},\dots,v_{l}\big{\}}$ is in bijection (the string parametrization) with the so-called dual crystal basis $\big{\{}e_{v_{1}}^{(n\lambda)},\cdots,e_{v_{l}}^{(n\lambda)}\big{\}}$ of $H^{0}(G/P,nM_{\lambda})$ (see [16, §3] for a detailed exposition). Moreover, the dual crystal basis is compatible with the weight decomposition
$H^{0}(G/P,nM_{\lambda})=\bigoplus_{\mu\in X(T)}H^{0}(G/P,nM_{\lambda})[\mu],$
that is, for each $i=1,\dots,l$, $e_{v_{i}}^{(n\lambda)}\in H^{0}(G/P,nM_{\lambda})[\mu]$ for some $\mu$.
The string polytope $\Delta_{\lambda}$ is the Okounkov body $\Delta_{\nu}(G/P,M_{\lambda})$, i.e. it is the closure of the set $\bigcup_{n\geq 1}\tfrac{1}{n}\nu(H^{0}(G/P,nM_{\lambda})\backslash\{0\})$. Note that $\Delta_{\lambda}\cap\tfrac{1}{n}\mathbb{Z}^{N}$ is in bijection with the dual crystal basis of $H^{0}(G/P,nM_{\lambda})$.
The weight polytope $\mathcal{P}_{\lambda}\subseteq X(T)_{\mathbb{R}}$ is the convex hull of the Weyl orbit $W\lambda$ in $X(T)_{\mathbb{R}}$. Note that $\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)$ is in bijection with weights apprearing in $H^{0}(G/P,nM_{\lambda})$. By taking weights, one has weight map $p:\Delta_{\lambda}\longrightarrow\mathcal{P}_{\lambda}$ which is affine and sends $\Delta_{\lambda}\cap\tfrac{1}{n}\mathbb{Z}^{N}$ to $\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)$
3.3. Proof of Theorem 3.1
Let $V_{n}=H^{0}(G/P,nM_{\lambda})$ and $V_{\bullet}=\bigoplus_{n\geq 0}V_{n}$. For $t\in\mathbb{R}$, let $V_{n}^{t}=H^{0}(G/P,nM_{\lambda})_{t,{\mathrm{deg}}(F_{Q})}$ as in Proposition 2.12, and $V_{\bullet}^{t}=\bigoplus_{n\geq 0}V_{n}^{nt}$. Let $\nu:V_{\bullet}\backslash\{0\}\longrightarrow\mathbb{Z}^{N}$ be the string valuation. Let $\varphi:\Delta_{\nu}(G/P,M_{\lambda})=\Delta_{\lambda}\longrightarrow\mathbb{R}$ be the Boucksom-Chen concave transform associated to $\left(V_{\bullet},\nu,\big{\{}V_{\bullet}^{t}\big{\}}\right)$.
Theorem 3.3.
$\varphi(x)=\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle$.
@@ 3.2. String polytopes and weight polytopes
$V_{n}^{nt}=\bigoplus_{\mu}V_{n}[n\mu]$ for $\mu\in\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)$ with $\langle{\mathrm{deg}}(F_{Q}),n\mu\rangle\geq nt$, and hence $V_{n}^{nt}=\bigoplus_{x}k\cdot e_{nx}^{(n\lambda)}$ for $x\in\Delta_{\lambda}\cap\tfrac{1}{n}\mathbb{Z}^{N}$ with $\langle{\mathrm{deg}}(F_{Q}),p(nx)\rangle\geq nt$. Therefore by definition,
$\quad\quad\quad\Delta(V_{\bullet}^{t})$ $=\text{closure of }\bigcup\nolimits_{n}\big{\{}x\in\Delta_{\lambda}\cap\tfrac{1}{n}\mathbb{Z}^{N}:\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle\geq t\big{\}}$
=$=\big{\{}x\in\Delta_{\lambda}:\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle\geq t\big{\}}$
which concludes the proof. ∎
Consider the graded linear series $\widetilde{V}_{\bullet}=\bigoplus_{n\geq 0}H^{0}(X,nL)$. Fix an isomorphism $F_{Q,K}\simeq Q_{K}$. Then
(3.1) $\widetilde{V}_{\bullet}=\bigoplus_{n\geq 0}\Big{(}F_{Q}\times_{Q}H^{0}(G/P,nM_{\lambda})\Big{)}_{K}\simeq\bigoplus_{n\geq 1}H^{0}(G/P,nM_{\lambda})_{K}.$
Consider the filtrations of $\widetilde{V}_{n}=\Big{(}F_{Q}\times_{Q}H^{0}(G/P,nM_{\lambda})\Big{)}_{K}$ induced by the Harder-Narasimhan filtration of $\pi_{*}\mathcal{L}_{n\lambda}=F_{Q}\times_{Q}H^{0}(G/P,nM_{\lambda})$ as in Proposition 2.12
$\widetilde{V}_{\bullet}^{t}=\bigoplus_{n\geq 0}\Big{(}F_{Q}\times_{Q}H^{0}(G/P,nM_{\lambda})_{nt,{\mathrm{deg}}(F_{Q})}\Big{)}_{K}\simeq\bigoplus_{n\geq 0}\Big{(}H^{0}(G/P,nM_{\lambda})_{nt,{\mathrm{deg}}(F_{Q})}\Big{)}_{K}.$
Via the string parametrization of the dual crystal basis, the string valuation $\nu$ on $H^{0}(G/P,nM_{\lambda})$ extends canonically to a valuation $\widetilde{\nu}$ on $H^{0}(G/P,nM_{\lambda})_{K}$. This induces a valuation on $\widetilde{V}_{n}\simeq H^{0}(G/P,nM_{\lambda})_{K}$ via the isomorphism (3.1). We shall abuse notation and also call it $\widetilde{\nu}$.
The Boucksom-Chen concave transform of ${\mathcal{L}_{\lambda}}$ with respect to $\widetilde{\nu}$ is the one associated to $\left(\widetilde{V}_{\bullet},\widetilde{\nu},\big{\{}\widetilde{V}_{\bullet}^{t}\big{\}}\right)$, and obviously it coincides with the one associated to $\left(V_{\bullet},\nu,\big{\{}V_{\bullet}^{t}\big{\}}\right)$. Theorem 3.1 follows.
3.4. Proof of Theorem 3.2
@@
Let $\mu$ be the Lebesgue measure on the real span of $\Delta_{\lambda}$, normalized with respect to $\mathbb{Z}^{N}$. By [6], we have
@@
$h_{\mathcal{L}_{\lambda}}(X)=\frac{\mathcal{L}_{\lambda}^{\dim\mathcal{X}}}{\dim\mathcal{X}\cdot L^{\dim X}}=\frac{1}{\operatorname{Vol}(\Delta_{\lambda})}\int_{\Delta_{\lambda}}\langle{\mathrm{deg}}(F_{Q}),p(x)\rangle\,\mathrm{d}\mu(x).$
@@
For any linear function $f:X(T)_{\mathbb{R}}\longrightarrow\mathbb{R}$, we have
@@
Proof of Theorem 3.2.
Take $f={\mathrm{deg}}(F_{Q})$ in Proposition 3.4. ∎
The rest of this subsection is devoted to the proof of Proposition 3.4. We start with two lemmata.
@@
Note that the measure $\mu$ on $\Delta_{\lambda}$ is the limit of the Dirac measures $\delta_{\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}$, i.e. $\mu=\lim_{n}\delta_{\Delta_{\lambda}\cap\frac{1}{n}\mathbb{Z}^{N}}$. Consider the map $\Delta_{n\lambda}\longrightarrow\mathcal{P}_{n\lambda}$ and note that it sends integral points to its corresponding weights. Dividing by n, we see that
@@
$\operatorname{mult}_{n}:\mathcal{P}_{\lambda}\cap\tfrac{1}{n}X(T)\longrightarrow\mathbb{N},\quad\mu\longmapsto\dim H^{0}(G/P,M_{n\lambda})[n\mu].$
@@
$H^{0}(G/P,M_{\lambda})$ is a simple $G$-module so all weights $\mu$ occuring in $H^{0}(G/P,M_{\lambda})$ share a same restriction to $Z(G)$. So by the above claim, the lemma is true for lattice points $\mathcal{P}_{\lambda}\cap X(T)$. Note that the map $\mu\longmapsto\sum_{w\in W}w\mu$ is linear, we are done. ∎
@@
We end this section by the case $\operatorname{Gr}(2,4)$.